OpenStudy (anonymous):

What is the similarity ratio of the smaller to larger similar cones? larger one v=1024 pi m^3 smaller one v=250 pi m^3

OpenStudy (anonymous):

Use the formula V1 / V2 = (l1 / l2)^3 l1 : l2 is the similarity ratio

OpenStudy (anonymous):

does that say 11:12. just want to make sure

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Length 1 and length 2

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

(1024/250)^3=68.71